Ammonium nitrate has been used as a high explosive because it is unstable and decomposes into several gaseous substances. The rapid expansion of the gaseous substances produces the explosive force. NH4NO3 (s) --> N2 (g) + O2 (g) + H20 (g) Calculate the mass of each product gas if 20.0 g of ammonium nitrate reacts.
Answer #1

The molecular weight of ammonium nitrated is 80 g/mole.

The balance chemical equation of ammonium nitrate :

2NH4NO3 ---> 2N2 + O2 +4H2O

So 160 grams of ammonium nitrate produces 56 g Nitrogen, 32 gram Oxygen and 72 g of water.

so if 20 grams of ammonium nitrate ie 1/8 of 160 g of reactant ammonium nitrate taken than the product  form will also be 1/8 .

160 x 1/8 = 20 gram ammonium nitrate taken.

Nitrogen produced 56 x 1/8 = 7 g

Oxygen produced 32 x 1/8 = 4 g

Water produced 72 x 1/8 = 9 g.

Mass of product N2 7 g, O2 4g and H2O 9 g will be produce by 20 g NH4NO3.